# IndicGenBench: A Multilingual Benchmark to Evaluate Generation Capabilities of LLMs on Indic Languages

### v1

@@ Abstract
As large language models (LLMs) see increasing adoption across the globe, it is imperative for LLMs to be representative of the linguistic diversity of the world. India is a linguistically diverse country of 1.4 Billion people. To facilitate research on multilingual LLM evaluation, we release IndicGenBench — the largest benchmark for evaluating LLMs on user-facing generation tasks across a diverse set 29 of Indic languages covering 13 scripts and 4 language families. IndicGenBench is composed of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering. IndicGenBench extends existing benchmarks to many Indic languages through human curation providing multi-way parallel evaluation data for many under-represented Indic languages for the first time. We evaluate a wide range of proprietary and open-source LLMs including GPT-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM and LLaMA on IndicGenBench in a variety of settings. The largest PaLM-2 models performs the best on most tasks, however, there is a significant performance gap in all languages compared to English showing that further research is needed for the development of more inclusive multilingual language models. IndicGenBench is available at www.github.com/google-research-datasets/indic-gen-bench
@@ 1 Introduction
We release IndicGenBench, a multilingual, multi-way parallel benchmark for measuring language generation capabilities across diverse user-facing tasks in 29 Indic languages across 4 language families (Table 7). IndicGenBench extends existing benchmarks such as CrossSum (Bhattacharjee et al., 2023), XQuAD (Artetxe et al., 2020), XorQA (Asai et al., 2021), and FLORES (NLLB-Team et al., 2022) for additional Indic languages and is composed of tasks like cross-lingual summarization (CrossSum-In), machine translation (Flores-In), cross-lingual reading comprehension (XorQA-In-Xx and XorQA-In-En) and multilingual reading comprehension (XQuAD-In). Each dataset consists of parallel examples in up to 29 low to comparatively higher resource Indic languages; and for some tasks (e.g. CrossSum-In), IndicGenBench provides the first-ever evaluation datasets for as many as 18 of these languages. We also release a small training set in all tasks for efficient adaptation of LLMs. Our comprehensive evaluation of various state-of-the-art proprietary and open-source LLMs on IndicGenBench shows that there is a significant gap in performance between English and Indic languages (see Figure 1). Our contributions are as follows:
@@ 2 IndicGenBench
1. XorQA-In-En: Each example contains (Xx-question, En-passage, En-answer). This task is similar to the Xor-TyDi dataset in additional Indic languages.
@@ 3 Experiments and Analysis
We compare and analyze the performance of these LLMs and their variants in terms of model sizes under different learning paradigms and settings. We first evaluate model performance on one-shot prompting (§3.1) and also measure performance across language categories based on resourcedness (§3.2). We then evaluate the effect of number of in-context examples shown to the model as supervised data (§3.3) and the effect of prompting in a higher-resource language such an English or Hindi (§3.4). Using the training data contained in IndicGenBench, we measure how the performance of LLMs after fine-tuning compares with few-shot prompting (§3.5). Finally, we perform qualitative analysis of models on IndicGenBench and highlight some areas of improvement for future model development (§3.7).
@@ 3.2 Performance across language categories
(a) In Flores-In, the performance for translating English to the target language (enxx) drops significantly from higher to lower resourced languages ( for PaLM-2-L) whereas the performance in the xxen direction does not fall this drastically (). A similar trend is seen when comparing XorQA-In-Xx and XorQA-In-En. This highlights that current LLMs are better at understanding these lower-resourced languages than generating fluent text in them.
@@ 3.6 Analyzing Tokenizer across Indic languages
A high token fertility is undesirable and can disproportionately effect a particular language’s performance. For languages where text is broken into more number of tokens, fewer in-context examples can be input to the LLM during inference. This can negatively impact performance (see Table 4). In Figure 3, we show how the percentage of data that fits in a particular context length changes with number of in-context examples for various languages. For example, we see in Figure 3 that for medium resource languages with high token-fertility like Oriya and Punjabi we can in-corporate much fewer in-context examples, compared to Indic languages with lower token-fertility like Hindi and Marathi.
@@ 4 Related Work
In the last few years, many multilingual LLMs have been developed—starting from mBART (Liu et al., 2020) trained on 25 languages to LLMs that are pre-trained on hundreds of languages, such as mT55 (Xue et al., 2021), PaLM-2 (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Gemini (Google, 2023), and others. These LLMs are typically evaluated on individual multilingual tasks for Translation: WMT (Farhad et al., 2021), FLORES (NLLB-Team et al., 2022); Question-Answering: XQuAD (Artetxe et al., 2020), TyDiQA (Clark et al., 2020), XorQA (Asai et al., 2021); Summarization: XLSUM (Hasan et al., 2021a), CrossSum (Bhattacharjee et al., 2023); Reasoning: MGSM (Shi et al., 2022), XCOPA (Ponti et al., 2020) to name a few, or on multilingual benchmarks such as, XTREME (Hu et al., 2020) and XTREME-UP (Ruder et al., 2023). However, most of these evaluation resources contain only a handful of languages or do not contain data for low resource languages, especially Indics. Besides, cross-lingual evaluation data is even more sparse. This work is an effort to bridge these gaps by releasing IndicGenBench, a suite of datasets covering diverse cross-lingual and multilingual generation tasks in Indic languages.
@@ Appendix E Dataset Licenses
In compliance with licenses of the original datasets, we release our evaluation datasets under the following licenses: (a) XQuAD-In andFlores-In under the CC-BY-SA 4.0 license; (b) CrossSum-In under the CC BY-NC-SA 4.0 license; and (d) XorQA-In under the MIT license.

### v2

@@ Abstract
As large language models (LLMs) see increasing adoption across the globe, it is imperative for LLMs to be representative of the linguistic diversity of the world. India is a linguistically diverse country of 1.4 Billion people. To facilitate research on multilingual LLM evaluation, we release IndicGenBench — the largest benchmark for evaluating LLMs on user-facing generation tasks across a diverse set 29 of Indic languages covering 13 scripts and 4 language families. IndicGenBench is composed of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering. IndicGenBench extends existing benchmarks to many Indic languages through human curation providing multi-way parallel evaluation data for many under-represented Indic languages for the first time. We evaluate a wide range of proprietary and open-source LLMs including GPT-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM and LLaMA on IndicGenBench in a variety of settings. The largest PaLM-2 models performs the best on most tasks, however, there is a significant performance gap in all languages compared to English showing that further research is needed for the development of more inclusive multilingual language models. IndicGenBench is available at www.github.com/google-research-datasets/indic-gen-bench ††All authors are now part of Google DeepMind
@@ 1 Introduction
We release IndicGenBench, a multilingual, multi-way parallel benchmark for measuring language generation capabilities across diverse user-facing tasks in 29 Indic languages across 4 language families (Table 7). IndicGenBench extends existing benchmarks such as CrossSum (Bhattacharjee et al., 2023), XQuAD (Artetxe et al., 2020), XorQA (Asai et al., 2021), and FLORES (NLLB-Team et al., 2022) for additional Indic languages and is composed of tasks like cross-lingual summarization (CrossSum-In), machine translation (Flores-In), cross-lingual reading comprehension (XorQA-In-Xx and XorQA-In-En) and multilingual reading comprehension (XQuAD-In). Each dataset consists of parallel examples in up to 29 low to comparatively higher resource Indic languages; and for some tasks (e.g. CrossSum-In), IndicGenBench provides the first-ever evaluation set for as many as 18 of these languages. We also release a small training set in all tasks for efficient adaptation of LLMs. Our comprehensive evaluation of various state-of-the-art proprietary and open-source LLMs on IndicGenBench shows that there is a significant gap in performance between English and Indic languages (see Figure 1). Our contributions are as follows:
@@ 2 IndicGenBench
1. XorQA-In-En: Each example contains (Xx-question, En-passage, En-answer). This task is similar to the Xor-TyDi dataset.
@@
See Appendix Table 9 for languages covered by each dataset in IndicGenBench.
@@ 3 Experiments and Analysis
We compare and analyze the performance of different model size variants of these LLMs under various learning paradigm settings. We first evaluate model performance on one-shot prompting (§3.1) and also measure performance across language categories based on resourcedness (§3.2). We then evaluate the effect of number of in-context examples shown to the model as supervised data (§3.3) and the effect of prompting in a higher-resource language such an English or Hindi (§3.4). Using the training data contained in IndicGenBench, we measure how the performance of LLMs after fine-tuning compares with few-shot prompting (§3.5). Finally, we perform qualitative analysis of models on IndicGenBench and highlight some areas of improvement for future model development (§3.7).
@@ 3.2 Performance across language categories
(a) In Flores-In, the performance for translating English to the target language (enxx) drops significantly from higher to lower resourced languages ( for PaLM-2-L) whereas the performance in the xxen direction does not fall this drastically (). This is also seen in XorQA-In-Xx and XorQA-In-En. This highlights that current LLMs are better at understanding than generation in these lower-resourced languages.
@@ 3.6 Analyzing Tokenizer across Indic languages
A high token fertility is undesirable and can disproportionately effect a particular language’s performance. For languages where text is broken into more number of tokens, fewer in-context examples can be input to the LLM during inference. This can negatively impact performance (see Table 4). In Figure 3, we show how the percentage of a dataset that fits in a particular context length changes with number of in-context examples for various languages. For example, we see in Figure 3 that for medium resource languages with high token-fertility like Oriya and Punjabi we can in-corporate much fewer in-context examples compared to Indic languages with lower token-fertility like Hindi and Marathi. Analysis in this section is inspired by prior work on studying the impact of tokenization on model utility, inference, and financial cost (Ahia et al., 2023; Petrov et al., 2023). Our work supports and compliments these studies for a range of Indic languages. Alternate tokenization and encoding approaches are required to bridge the tokenization disparity between different languages. An example of this is MYTE (Limisiewicz et al., 2024), a morphology based byte encoding scheme which leads to more equitable text representations.
@@ 4 Related Work
In the last few years, many multilingual LLMs have been developed—starting from mBART (Liu et al., 2020) trained on 25 languages to LLMs that are pre-trained on hundreds of languages, such as mT5 (Xue et al., 2021), PaLM-2 (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Gemini (Google, 2023), and others. These LLMs are typically evaluated on individual multilingual tasks for Translation: WMT (Farhad et al., 2021), FLORES (NLLB-Team et al., 2022); Question-Answering: XQuAD (Artetxe et al., 2020), TyDiQA (Clark et al., 2020), XorQA (Asai et al., 2021); Summarization: XLSUM (Hasan et al., 2021a); Reasoning: MGSM (Shi et al., 2022), XCOPA (Ponti et al., 2020) to name a few, or on multilingual benchmarks such as, XTREME (Hu et al., 2020) and XTREME-UP (Ruder et al., 2023). However, most of these evaluation resources contain only a handful of languages or do not contain data for low resource languages, especially Indics. Besides, cross-lingual evaluation data is even more sparse. This work is an effort to bridge these gaps by releasing IndicGenBench, a suite of datasets covering diverse cross-lingual and multilingual generation tasks in Indic languages.
@@ B.5 Indic Specific LLMs
| Code | Language | Coverage in IndicGenBench | ||||
| CrossSum-In | Flores-In | XQuAD-In | XorQA-In-Xx | XorQA-In-En | ||
| bn | Bengali | ✓ | ✓ | ✓ | ✓ | ✓ |
| gu | Gujarati | ✓ | ✓ | ✓ | ✓ | ✓ |
| hi | Hindi | ✓ | ✓ | ✓ | ✓ | ✓ |
| kn | Kannada | ✓ | ✓ | ✓ | ✓ | ✓ |
| ml | Malayalam | ✓ | ✓ | ✓ | ✓ | ✓ |
| mr | Marathi | ✓ | ✓ | ✓ | ✓ | ✓ |
| ta | Tamil | ✓ | ✓ | ✓ | ✓ | ✓ |
| te | Telugu | ✓ | ✓ | ✓ | ✓ | ✓ |
| ur | Urdu | ✓ | ✓ | ✓ | ✓ | ✓ |
| as | Assamese | ✓ | ✓ | ✓ | ✓ | ✓ |
| bho | Bhojpuri | ✓ | ✓ | ✓ | ✓ | |
| ne | Nepali | ✓ | ✓ | |||
| or | Odia | ✓ | ✓ | ✓ | ✓ | ✓ |
| pa | Punjabi | ✓ | ✓ | ✓ | ✓ | ✓ |
| ps | Pashto | ✓ | ✓ | ✓ | ✓ | |
| sa | Sanskrit | ✓ | ✓ | ✓ | ✓ | |
| awa | Awadhi | ✓ | ✓ | ✓ | ✓ | |
| bgc | Haryanvi | ✓ | ✓ | ✓ | ✓ | |
| bo | Tibetan | ✓ | ✓ | ✓ | ✓ | |
| brx | Bodo | ✓ | ✓ | ✓ | ✓ | |
| gbm | Garhwali | ✓ | ✓ | ✓ | ✓ | |
| gom | Konkani | ✓ | ✓ | ✓ | ✓ | |
| hne | Chhattisgarhi | ✓ | ✓ | ✓ | ✓ | |
| hoj | Rajasthani | ✓ | ✓ | ✓ | ✓ | |
| mai | Maithili | ✓ | ✓ | ✓ | ✓ | |
| mni | Manipuri | ✓ | ✓ | ✓ | ✓ | |
| mup | Malvi | ✓ | ✓ | ✓ | ✓ | |
| mwr | Marwari | ✓ | ✓ | ✓ | ✓ | |
| sat | Santali | ✓ | ✓ | ✓ | ✓ | |
@@ Appendix E Dataset Licenses
In compliance with licenses of the original datasets, we release our evaluation datasets under the following licenses: (a) XQuAD-In and Flores-In under the CC BY-SA 4.0 license; (b) CrossSum-In under the CC BY-NC-SA 4.0 license; and (d) XorQA-In under the MIT license.